OpenStudy (kanwal32):

A galvanometer of resistance G is converted into a voltmeter to measure upto V volts by connecting a resistance R1 in series with the coil. If a resistance R2 is connected in series with it, then it can measure upto V/2 volts. Find the resistance, in terms of R1 and R2, required to be connected to convert it into a voltmeter that can read upto 2 V. Also find the resistance G of the galvanometer in terms of R1 and R2.

OpenStudy (kanwal32):

my g is coming G=R2-2R1

OpenStudy (kanwal32):

@Michele_Laino

OpenStudy (kanwal32):

@Kainui

OpenStudy (kanwal32):

@ganeshie8

OpenStudy (kropot72):

The current at full scale deflection is given by: \[\large I _{fsd}=\frac{V}{G+R _{1}}=\frac{\frac{V}{2}}{G+R _{2}}\] \[\large \frac{V}{G+R _{1}}=\frac{V}{2G+2R _{2}}\] \[\large G+R _{1}=2G+2R _{2}\] \[\large G=R _{1}-2R _{2}\]